A stock had returns of 18.58 percent, −5.58 percent, 20.81 percent, and 8.89 percent for the past four years. What is the variance of the returns?

  • .12009

  • .01442

  • .00648

  • .01923

  • .01730
Answer #1

Average return=Total return/Total time period

=(18.58−5.58+20.81+8.89)/4=10.675%

Return (Return-Average return)^2
18.58 (18.58-10.675)^2=62.489025
−5.58 (−5.58-10.675)^2=264.225025
20.81 (20.81-10.675)^2=102.718225
8.89 (8.89-10.675)^2=3.186225
Total=432.6185%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[432.6185/(4-1)]^(1/2)

=[432.6185/3]^(1/2)

=12.01%(Approx).

Variance=Standard deviation^2

=0.01442(Approx).